# HiKE: Hierarchical Evaluation Framework for Korean-English Code-Switching Speech Recognition

### v3

@@ 3 HiKE
SenseVoice-Small An et al. (2024) 234M 29.3 37.9 26.8 33.9 53.9 56.6 38.8 54.6 6.4 7.6
Whisper-Tiny Radford et al. (2023) 38M 78.8 128.7 39.1 103.8 89.1 76.4 68.5 81.2 11.9 14.6
Whisper-Base Radford et al. (2023) 74M 124.7 84.2 28.3 97.8 92.2 76.5 42.4 81.2 7.8 9.8
Whisper-Small Radford et al. (2023) 244M 59.2 46.0 20.6 50.1 69.3 46.7 31.6 55.2 4.5 8.3
Whisper-Medium Radford et al. (2023) 769M 44.6 33.6 17.9 37.3 40.6 51.7 31.1 46.1 3.4 4.6
Whisper-Large Radford et al. (2023) 1.5B 30.6 23.5 12.3 25.8 43.7 28.9 18.9 34.4 3.2 4.4
Seamless-M4T-v2-large Barrault et al. (2023) 2.3B 115.0 71.7 65.1 89.0 78.1 64.7 58.2 69.8 6.4 6.3
Gemma-3n Gemma Team (2025) 8B 61.6 64.0 57.7 62.7 68.5 65.9 48.4 66.0 10.7 13.0
Audio-Flamingo-3 Goel et al. (2025) 8.3B 79.2 83.7 80.8 81.7 102.2 98.2 116.0 100.7 25.1 8.8
GPT-4o-transcribe Hurst et al. (2024) N/A 17.8 27.1 27.5 23.3 25.9 30.8 33.5 28.9 2.5 3.3
@@ 3.5 Loanwords Post-processing
Loanwords are words adopted from a foreign language and adapted to the phonology and orthography of the new language. For example, the Korean loanword `버스' [bs] and the English word `bus' [bs] are pronounced almost identically and can be used interchangeably in a CS context. This creates an evaluation challenge: If a ground-truth label is strictly constrained to either Hangul or the Roman alphabet, a model producing the alternate—yet perfectly valid—transcription would be unfairly penalized. This ambiguity introduces significant noise into the assessment of CS performance. To avoid this problem, we meticulously labeled all loanwords contained in our dataset. Subsequently, during evaluation, both the Korean and English versions of these loanwords were treated as valid answers. Through our loanword labeling process, we achieved a more precise evaluation by decreasing measurement noise by an average of $5.1\%$ in MER and $5.9\%$ in PIER.
@@
Whisper-Medium 44.6 33.6 17.9 37.3 40.6 51.7 31.1 46.1 3.4 4.6
(a) FT with Natural Intra-Sentential CS Data 9.9(-34.7) 10.3(-23.3) 8.0(-9.9) 10.0(-27.3) 19.8(-20.8) 19.1(-32.6) 22.7(-8.4) 19.5(-26.6) 6.0(+2.6) 5.2(+0.6)
(b) FT with Synthetic Inter-sentential CS Data 21.0(-23.6) 27.6(-6.0) 7.5(-10.4) 23.9(-13.4) 33.0(-7.6) 32.9(-18.8) 16.0(-15.1) 32.1(-14.0) 3.1(+0.3) 5.7(+0.5)
(a) + (b) FT with Both Data 22.6(-22.0) 22.7(-10.9) 6.0(-11.9) 21.8(-15.5) 27.9(-12.7) 36.9(-14.8) 13.5(-17.6) 32.0(-14.1) 3.9(+0.5) 4.9(+0.3)
@@ 4.1 CS-ASR of Multilingual ASR Models
We evaluated $10$ multilingual ASR models with diverse architectures (CTC, Transformer, and LLM-based), assessing their CS-ASR performance using MER and PIER metrics. As shown in Table 2, a severe performance drop occurred on CS data across all models, with the MER increasing by a factor of $3$ to $14$ compared to monolingual performance, revealing significant limitations for practical use. A closer look at specific architectures reveals further nuances. For instance, the CTC-based SenseVoice-Small, despite showing stable performance across CS-levels and a better overall MER than the similarly-sized Whisper-Small, still exhibited a high error rate at the actual code-switching points. This suggests that while CTC architectures may be robust in overall transcription, they struggle specifically with the precise moment of language transition.
In contrast, models leveraging Large Language Models (LLMs) trained on extensive text corpora displayed distinct behavioral patterns compared to speech-centric models. Specifically, despite having over five times the parameter count of Whisper-Large, Gemma-3n and Audio-Flamingo-3 failed to achieve competitive results in CS-ASR, with both MER and PIER exceeding $60$. GPT-4o-Transcribe was the only LLM-based model to outperform Whisper-Large. Notably, while most non-LLM models performed best on sentence-level CS and worst on word-level CS, GPT-4o-Transcribe exhibited the opposite trend, achieving its highest performance on word-level CS and its weakest on sentence-level CS. We hypothesize that this reflects the distribution of its training data, as large text corpora are rich in word-level CS but contain comparatively few instances of sentence-level CS.
To analyze the effect of model scale on CS-ASR performance, we evaluated the Whisper family of models. The results show a clear correlation between size and capability: CS-ASR performance, nearly absent in the smallest models (e.g., Tiny and Base), gradually emerges with increasing scale. Despite this trend, even the largest model's error rate on CS data was over six times higher than on monolingual data, indicating that model scaling alone is an insufficient solution for achieving practical CS-ASR performance.
@@ 4.2.2 Results
The results in Table 3 reveal that fine-tuning is an effective method for enabling CS-ASR, and that this can be achieved not only with natural intra-sentential CS data (a) but also with synthetic inter-sentential CS data (b) created by concatenating monolingual utterances. As shown in row (b) of Table 3, fine-tuning of synthetic CS data alone improved both the overall MER and PIER by more than $13\%$. Given the wide availability and relative ease of collecting monolingual data compared to authentic CS data, this finding suggests that data synthesis via utterance concatenation represents a promising and cost-effective direction for training CS-ASR models, especially in resource-constrained scenarios. However, fine-tuning on natural intra-sentential CS data (a) yielded greater performance improvements than fine-tuning on synthetic inter-sentential data (b, c) across almost all metrics; the only exceptions were the MER and PIER for sentence-level CS. Although this result indicates that it is currently preferable to use natural intra-sentential CS data for fine-tuning when available, we anticipate that these limitations can be overcome. With the development of more sophisticated data synthesis or fine-tuning techniques, it may become possible to train robust CS-ASR models using only synthetic data.
@@ Acknowledgments
We appreciate Hyunwoo Kim for proofreading and the anonymous reviewers for their insightful feedback and suggestions. We also thank the Theta One Team for supporting our work.

### v4

@@ 3 HiKE
SenseVoice-Small An et al. (2024) 234M 27.4 36.4 23.6 32.5 52.2 56.5 37.1 54.7 6.4 7.6
Whisper-Tiny Radford et al. (2023) 38M 73.7 111.4 36.2 93.6 81.7 77.9 66.9 78.9 11.9 14.6
Whisper-Base Radford et al. (2023) 74M 116.5 81.8 25.4 91.2 90.2 73.2 38.7 78.1 7.8 9.8
Whisper-Small Radford et al. (2023) 244M 50.6 41.3 19.9 43.5 58.0 46.7 31.5 50.1 4.5 8.3
Whisper-Medium Radford et al. (2023) 769M 39.1 27.9 16.8 31.3 41.9 41.3 30.6 41.3 3.4 4.6
Whisper-Large Radford et al. (2023) 1.5B 28.5 25.3 20.0 26.1 45.6 31.3 25.8 36.0 3.2 4.4
Seamless-M4T-v2-large Barrault et al. (2023) 2.3B 108.8 70.1 61.0 83.6 72.4 60.9 57.3 64.7 6.4 6.3
Gemma-3n Gemma Team (2025) 8B 99.6 64.6 54.3 76.6 69.2 47.8 40.3 54.8 10.7 13.0
Audio-Flamingo-3 Goel et al. (2025) 8.3B 78.5 82.2 79.4 80.7 104.1 97.8 112.9 100.2 25.1 8.8
GPT-4o-transcribe Hurst et al. (2024) N/A 15.6 25.0 28.3 21.8 25.8 30.2 32.3 28.8 2.5 3.3
@@ 3.5 Loanwords Post-processing
Loanwords are words adopted from a foreign language and adapted to the phonology and orthography of the new language. For example, the Korean loanword `버스' [bs] and the English word `bus' [bs] are pronounced almost identically and can be used interchangeably in a CS context. This creates an evaluation challenge: If a ground-truth label is strictly constrained to either Hangul or the Roman alphabet, a model producing the alternate—yet perfectly valid—transcription would be unfairly penalized. This ambiguity introduces significant noise into the assessment of CS performance. To avoid this problem, we meticulously labeled all loanwords contained in our dataset. Subsequently, during evaluation, both the Korean and English versions of these loanwords were treated as valid answers. Through our loanword labeling process, we achieved a more precise evaluation by decreasing measurement noise by an average of $4.9\%$ in MER and $7.9\%$ in PIER.
@@
Whisper-Medium 39.1 27.9 16.8 31.3 41.9 41.3 30.6 41.3 3.4 4.6
(a) FT with Natural Intra-Sentential CS Data 8.3(-30.8) 9.6(-18.4) 7.4(-9.4) 9.0(-22.3) 18.6(-23.3) 19.9(-21.4) 21.0(-9.7) 19.5(-21.8) 6.0(+2.6) 5.2(+0.6)
(b) FT with Synthetic Inter-sentential CS Data 19.1(-20.0) 25.6(-2.3) 5.8(-11.0) 22.1(-9.2) 33.7(-8.2) 35.6(-5.7) 14.5(-16.1) 34.5(-6.8) 3.7(+0.3) 5.1(+0.5)
(a) + (b) FT with Both Data 20.7(-18.4) 21.7(-6.3) 4.8(-11.9) 20.4(-11.0) 27.3(-14.6) 37.6(-3.7) 11.3(-19.4) 33.6(-7.7) 3.9(+0.5) 4.9(+0.3)
@@ 4.1 CS-ASR of Multilingual ASR Models
We evaluated $10$ multilingual ASR models with diverse architectures (CTC, Transformer, and LLM-based), assessing their CS-ASR performance using MER and PIER metrics. As shown in Table 2, a severe performance drop occurred on CS data across all models, with the MER increasing by a factor of $3$ to $13$ compared to monolingual performance, revealing significant limitations for practical use. A closer look at specific architectures reveals further nuances. For instance, the CTC-based SenseVoice-Small, despite showing stable performance across CS-levels and a better overall MER than the similarly-sized Whisper-Small, still exhibited a high error rate at the actual code-switching points. This suggests that while CTC architectures may be robust in overall transcription, they struggle specifically with the precise moment of language transition.
In contrast, models leveraging Large Language Models (LLMs) trained on extensive text corpora displayed distinct behavioral patterns compared to speech-centric models. Specifically, despite having over five times the parameter count of Whisper-Large, Gemma-3n and Audio-Flamingo-3 failed to achieve competitive results in CS-ASR, with both MER and PIER exceeding $54$. GPT-4o-Transcribe was the only LLM-based model to outperform Whisper-Large. Notably, while most non-LLM models performed best on sentence-level CS and worst on word-level CS, GPT-4o-Transcribe exhibited the opposite trend, achieving its highest performance on word-level CS and its weakest on sentence-level CS. We hypothesize that this reflects the distribution of its training data, as large text corpora are rich in word-level CS but contain comparatively few instances of sentence-level CS.
To analyze the effect of model scale on CS-ASR performance, we evaluated the Whisper family of models. The results show a clear correlation between size and capability: CS-ASR performance, nearly absent in the smallest models (e.g., Tiny and Base), gradually emerges with increasing scale. Despite this trend, even the largest model's error rate on CS data was over five times higher than on monolingual data, indicating that model scaling alone is an insufficient solution for achieving practical CS-ASR performance.
@@ 4.2.2 Results
The results in Table 3 reveal that fine-tuning is an effective method for enabling CS-ASR, and that this can be achieved not only with natural intra-sentential CS data (a) but also with synthetic inter-sentential CS data (b) created by concatenating monolingual utterances. As shown in row (b) of Table 3, fine-tuning of synthetic CS data alone improved both the overall MER and PIER by more than $6.8\%$. Given the wide availability and relative ease of collecting monolingual data compared to authentic CS data, this finding suggests that data synthesis via utterance concatenation represents a promising and cost-effective direction for training CS-ASR models, especially in resource-constrained scenarios. However, fine-tuning on natural intra-sentential CS data (a) yielded greater performance improvements than fine-tuning on synthetic inter-sentential data (b, c) across almost all metrics; the only exceptions were the MER and PIER for sentence-level CS. Although this result indicates that it is currently preferable to use natural intra-sentential CS data for fine-tuning when available, we anticipate that these limitations can be overcome. With the development of more sophisticated data synthesis or fine-tuning techniques, it may become possible to train robust CS-ASR models using only synthetic data.
@@ Acknowledgments
We appreciate Hyunwoo Kim for proofreading and the anonymous reviewers for their insightful feedback and suggestions. Special thanks to DongChan Shin (@DongChanS) for pointing out issues in the preprocessing pipeline on GitHub. We also thank the Theta One Team for supporting our work.